OpenStudy (emmynimmy):

Solve the equation for x. 12 − 5(x + 3) = 4x − 5

OpenStudy (hyuna301):

12 - 5x - 15 = 4x - 5 12 - 15 + 5 = 4x + 5x 2 = 9x x = 2/9

OpenStudy (hyuna301):

we can also write it as (4x - 9)/12 = (3x + 5)/12 4x - 9 = 3x + 5 x = 5 + 9 x = 14

OpenStudy (emmynimmy):

what about the second one? @Hyuna301

OpenStudy (hyuna301):

[ x + 5 - 2( x - 1 ) ] /6 = 5 x + 5 - 2x + 2 = 30 -x + 7 = 30 -x = 23 x = -23

OpenStudy (emmynimmy):

A wireless carrier offers the following plans that a person is considering. The Family Plan: $100 monthly fee, unlimited talk and text on up to 5 lines, and data charges of $30 for each device for up to 2 GB of data per device. The Mobile Share Plan: $160 monthly fee for up to 10 devices, unlimited talk and text for all the lines, and data charges of $20 for each device up to a shared total of 10 GB of data. Find the model of the total cost of the Family Plan in dollars. Use P for the number of devices that need data plans as part of their cost.

OpenStudy (psycho45):

I keep getting 3.5 no matter what I did